OpenStudy (anonymous):

In Triangle ABC, XY=6.32, YZ=5.83 and XZ=11.05. What is m

OpenStudy (solomonzelman):

I have imagination problems, and I am not very creative, can you post a picture please ?

OpenStudy (anonymous):

sure, it may not be to scale

OpenStudy (solomonzelman):

sure :)

OpenStudy (anonymous):



OpenStudy (anonymous):

Is the way I am imagining it to be

OpenStudy (anonymous):

do i use Sine or Cosine?

OpenStudy (solomonzelman):

Ohh I forgot the formula for this -:(

OpenStudy (anonymous):

i know the formulas if you need them

OpenStudy (solomonzelman):

Yes, what is the formula? Or any hint?

OpenStudy (anonymous):

the formula for Sine is \[\frac{ SinA }{ a }=\frac{ SinB }{ b }=\frac{ SinC }{ C }\]

OpenStudy (anonymous):

the c under SinC is lowercase

OpenStudy (anonymous):

Cosine is \[a^@=b^2+c^2-2bc*\cos(A)\]

OpenStudy (anonymous):

\[b^2=a^2+c^2-2ac*\cos(B)\]

OpenStudy (anonymous):

\[c^2=a^2+b^2-2ab*\cos(C)\]